What is the speed of sound in air if its temperature is 560°C
1
Expert's answer
2020-09-08T09:11:43-0400

The formula for the speed of sound in air is the following:


"c = 331m/s +0.6\dfrac{ m}{s\cdot \degree C}\cdot T"

where "T = 560\degree C".

Thus, obtain:


"c = 331m/s +0.6\dfrac{ m}{s\cdot \degree C}\cdot 560\degree C = 667 m/s"

Answer. 667 m/s.